OpenStudy (anonymous):

A game has 40 chips, each labeled with a number from 1 through 40. Ben randomly chooses one chip. What is the probability that he chooses 10 or a multiple of 8?

OpenStudy (anonymous):

A. 1/8 B. 3/20 C.1/6 D. 1/320

OpenStudy (tkhunny):

10, 8, 16, 24, 32, 40 ??

OpenStudy (anonymous):

So how many multiples of 8 are between 1 and 40. Add that number with 1 (because of the one chip that's 10). And divide that by 40 because of the 40 chips.

OpenStudy (anonymous):

5

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

Its B :D

OpenStudy (anonymous):

5 multiples plus 1 divided by 40 = .15= 3/20= 15/100

OpenStudy (anonymous):

Yep! Your right!